A town has a population of 12000 and grows at 3.5% every year. What will be the population after 7 years, to the nearest whole number?


1
Expert's answer
2021-03-12T01:16:10-0500

Today town has a population of 12000 that is 100%. After 1 year it will be 103.5%, i.e. Y1=12000*1.035 .

After 2 years: Y2=Y1*1.035=(12000*1.035)*1.035=12000*1.0352 .

In this way after 7 years population will be 12000*1.0357=15267.35

Answer (to the nearest whole number): 15267